Without using finance calculator,
(a) What sum of money will become RM 15,000 in three years at a simple interest rate of 8% per annum?
(b) A certain sum of money is invested now. This investment will be worth RM5500 after fifteen months and RM5800 after twenty-four months. Find the original principal and the simple interest rate that was offered.
1
Expert's answer
2020-05-27T09:18:28-0400

a) If the investment is x after 3 years it will become:

"x*(1+0.08*3)=15000"

Which means that sum of money is

"x=\frac{15000}{1.24}\approx12096.77" RM


b) If the initial investment is x and interest rate is r, investment after fifteen and twenty four month will be equial to:

"x*(1+\frac{15}{12}*r)=5500\\\\x*(1+\frac{24}{12}*r)=5800"


Which means that original principal is RM 5 000, offered simple interest rate is 8%.